OpenStudy (anonymous):

Tony had 3 1/2 cups of flour and used 1 1/5 cups for a cake. How many cups of flour did Tony have left?

OpenStudy (anonymous):

got 2 1/10 cups

OpenStudy (texaschic101):

3 1/2 - 1 1/5 (change to improper fractions) 7/2 - 6/5 (find common denominator which is 10) 35/10 - 12/10 = 23/10 = 2 3/10 cups < -- flour that is left

OpenStudy (mathstudent55):

Method 1: Convert mixed numerals to fractions and subtract. Then convert result back to mixed numeral.